a+8b=8
a+3b=-2
Use the elimination method to solve system of equations. Thanks

Found 2 solutions by mananth, waynest:
Answer by mananth(16946)

1 a + 8 b = 8 .............1
Total value
1 a + 3 b = -2 .............2
Eliminate b
subtract (2) from (1)
5b=10
b=2
substitute b in (1)
a+8b=8
a+8*2=8
a+16=8
a=-16+8
a=-8


Answer by waynest(281)

a + 8b = 8
-(a + 3b = -2)
5b = 10
b = 2
a + 8(2) = 8
a + 16 = 8
a = 8 -16
a = -8
check:
-8 + 8(2) = 8
-8 + 16 = 8
8 = 8
-8 + 3(2) = -2
-8 + 6 = -2
-2 = -2